One year consumers spent an average of ​$23 on a meal at a resturant. Assume that the amount spent on a resturant meal is normally distributed and that the standard deviation is

​$5.

Complete parts​ (a)

a. What is the probability that a randomly selected person spent more than ​$25​?
Answer #1

Part a)

X ~ N ( µ = 23 , σ = 5 )
P ( X > 25 ) = 1 - P ( X < 25 )
Standardizing the value
Z = ( X - µ ) / σ
Z = ( 25 - 23 ) / 5
Z = 0.4
P ( ( X - µ ) / σ ) > ( 25 - 23 ) / 5 )
P ( Z > 0.4 )
P ( X > 25 ) = 1 - P ( Z < 0.4 )
P ( X > 25 ) = 1 - 0.6554
P ( X > 25 ) = 0.3446